OpenStudy (anonymous):

what fraction of a centimeter equals 1.0 um (micrometer)? I'm not sure how to approach this, havent taken physics in over 4 years

OpenStudy (anonymous):

first, convert the micrometer to meter 1.0 um = 10^6 meter then convert meter to centimeter 1m = 100cm 10^6m = 10^8 cm